OpenStudy (anonymous):

The length of a rectangle is 3 inches more than its width, and its perimeter is 22 inches. Find the width of the rectangle 4 in. 5 in. 7 in.

Parth (parthkohli):

First - the perimeter of a rectangle is 2(l + w) Let's say that the width is x, then the length would be x + 3. Substitute and equate. \( \color{Black}{\Rightarrow 2(x + (x + 3)) = 22}\) Dividing both sides by 2 is a good start. \( \color{Black}{\Rightarrow x + (x + 3) = 11}\) combine the like terms. \( \color{Black}{\Rightarrow 2x + 3 = 11 }\) Can you do it now?

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

thx

Parth (parthkohli):

Yw